OpenStudy (timinthor):

Please help! i will medal if correct! Which equations are equations for the line shown? Choose all answers that are correct. (-3, 1) and (3,4) are graphed, i found the slope and it is 1/2 when simplified

OpenStudy (welshfella):

Slope is correct

OpenStudy (welshfella):

To find the equation use the point slope form Y-y1 =m(x-x1) Where m = slope and (X1,y1) is a point on the Lin)